# Confidential Serverless Computing

### v1

@@ 4.4 Design Principles and Primitives
#2: Decoupled guest OS architecture. Conventional CVMs include the, often bloated [63], guest OS in their TCB. Hacher decouples the guest OS from trusted operations and removes its boot and measurement process from the function invocation path, mirroring ARM TrustZone’s principles, i.e., secure/normal world separation [18, 112], as performed in prior works, such as TLR (Trusted Language Runtime) [130].
After trusted monitor initialization, Hacher boots the untrusted guest OS, totally isolated from function execution environments. This decoupling excludes the guest OS kernel from the TCB, significantly reducing the attack surface and ensuring the guest OS vulnerabilities do not compromise the security of serverless functions. The guest OS handles only auxiliary tasks such as receiving requests and forwarding function invocations to the trusted monitor, maintaining compatibility with existing serverless frameworks. This reduces startup times compared to CVM-based deployments where cold starts involve full guest OS initialization and measurement.
#3: Trusted process templates. Existing confidential serverless deployments must initialize each function instance from scratch. Startup optimizations (e.g., snapshotting) are impractical [63] due to memory confidentiality (e.g., encryption) requirements. To this end, Hacher introduces trusted processes as the core abstraction for executing serverless functions.
Building on this execution model, Hacher implements its differential attestation protocol, which incrementally builds cumulative trust chains by reusing zygote and function measurements, requiring only the measurement of mutable components (e.g., input/output) during invocation, thus reducing attestation latency for both standalone and chained functions.
#4: Dynamically loadable LibOS architecture. To minimize function startup latency and its TCB while supporting diverse workloads (e.g., Python, Node.js), Hacher employs a dynamically loadable LibOS architecture [161, 134], adopting concepts of unikernel-based systems [98, 121, 161]. However, instead of requiring developers to build custom LibOS images, Hacher dynamically loads lightweight LibOS-based runtimes tailored to specific functions at runtime.
The LibOS forms the core of Hacher’s trusted processes, providing them with essential runtime functionalities. By dynamically loading minimal runtimes rather than full OSes or containers, Hacher reduces initialization overhead. This design allows on-demand instantiation of lightweight, pre-measured templates and avoids redundant software measurement steps during function invocation, further lowering startup latency without increasing developer effort.
#5: Confidential networking architecture. CVM I/O stacks (e.g., VirtIO [114], vsock [43]) suffer from performance penalties due to data copies and VM exits, and high CPU utilization for cryptographic operations [90, 109, 60]. Hacher’s key insight is to take advantage of its high function density (§ 3.5) and provide a monitor-mediated, high-performance networking architecture for CVMs.
Hacher leverages function co-location and employs a data-centric I/O approach [156] to optimize function communication via secure data objects, building on insights from prior work (e.g., CAP-VM [132], Nephele [97], Pheromone [156]). Hacher’s high function density increases co-location opportunities, and thus, it can also benefit from data-centric schedulers [2, 80, 156] that effectively co-locate related functions.
In Hacher, data objects serve as secure endpoints for trustlets to exchange data without traversing conventional I/O stacks. The trusted monitor manages data object allocation and page table permissions, granting appropriate access to producers and consumers, and ensures secure data exchange while maintaining isolation between functions. This design eliminates unnecessary cryptographic operations and redundant data copies, reducing I/O overheads. For functions that cannot be co-located, Hacher falls back to standard CVM I/O, ensuring compatibility with existing serverless frameworks.
5 Design
5.1 Trusted Monitor
The trusted monitor is Hacher’s core component. It maintains strict isolation between each serverless function and the guest OS through hardware-enforced CVM partitioning and page table configurations. It operates at the highest privilege level of the CVM partitioning domains (e.g., VMPL-0).
Trusted process management. The trusted monitor manages the CVM resources and preserves the state of the guest OS and trusted processes using process descriptors. A process descriptor includes the values for general-purpose and control registers, address space information (page tables), process ID, CVM-partitioning privilege level, object information (§ 5.4), and process state. The trusted monitor is responsible for allocating resources and scheduling the guest OS and the trusted processes. Internally, the descriptor state is linked to one vCPU state of the CVM, and the trusted monitor uses a special instruction to switch the context (e.g., AP creation call in SEV-SNP [14]). By executing the guest OS and the trusted processes at lower CVM partitioning levels, the trusted monitor governs their execution.
Programming interface. As a base for its execution environment, Hacher adopts a LibOS architecture [22, 118, 30, 53, 89, 26]. This aligns well with the requirements of trusted processes, where the main code runs at the ring3, and offers users maximum flexibility through a common POSIX-compatible interface with a minimal resource footprint. Hacher tailors the LibOS architecture for serverless execution by allowing the dynamic loading of functions without requiring the recompilation of other components. Furthermore, the LibOS provides a data-centric I/O [156] API for inter-function communication (§ 5.2) and relies on function provider policies that include workflow information (e.g., function chaining), measurements of the involved zygotes and functions, and, optionally, remote storage encryption keys for functions requiring external storage access.
@@ 5.1 Trusted Monitor
Monitor system calls. A serverless framework, running within the guest, interacts with the trusted monitor through the defined monitor system calls (Table 2). These calls facilitate the execution of serverless functions, including the creation of zygotes/trustlets, function invocation, performing attestation, and loading the provider’s policy. Additionally, the trusted monitor defines an interface between trusted processes for the LibOS to request backend operations (§ 5.2).
@@
Attestation service. The trusted monitor includes an attestation service that generates an attestation report for remote users to verify the integrity of Hacher and its function execution. On function execution, the attestation service calculates the measurement of function components following Hacher’s differential attestation principles (§ 5.5) to optimize the measurement process by caching previously computed results. The root-of-trust of Hacher relies on CVM’s attestation mechanism [14, 72], whose attestation report includes the measurement of the initial guest state and memory. The attestation service integrates this report into its final result.
@@ 5.2 LibOS Architecture
Hacher’s LibOS [22, 118, 30, 53, 89, 26] integrates a function runtime as well as a shim layer and filesystem with a nested namespace for efficient function execution.
@@
Filesystem. The LibOS provides an in-memory file system backed by a zygote, at the cost of a bloated image. Removing non-commonly-used files reduces image size but compromises serverless function flexibility and user experience. As a solution, Hacher employs a nested namespace [144] that integrates an embedded with an external filesystem. From the application’s perspective, all files appear at the same mount point (e.g., ‘/’). Internally, the LibOS first checks the embedded filesystem; if a file is missing, it retrieves it from the guest OS’s filesystem via the monitor. To ensure file integrity, the shim layer verifies files from the guest OS using provided measurements specified in a manifest [66, 148] embedded in the zygote image and aborts operations in case of a mismatch.
Data-centric I/O. For inter-function communication, the LibOS provides a data object abstraction to the trusted processes. This abstraction allows a function to create, get, and update data objects via the API shown in Table 3. Unlike conventional CVM networking that requires traversing entire I/O stacks with multiple context switches and encryption overhead [29, 90, 109, 60], Hacher employs a shared memory-based local object store [128, 156, 84, 119, 117, 28], enabling faster communication. This is especially effective for Hacher as it avoids the costly CVM network overhead (§ 3.3). When the serverless orchestrator fails to co-locate functions, the trusted monitor follows a hybrid I/O approach [99, 156] and resorts to the standard CVM I/O primitives via the guest OS, similar to the external file system access described above. Importantly, this mechanism ensures transparent communication between functions, regardless of whether they are co-located or distributed. § 5.4 details Hacher’s communication mechanism.
@@ 5.3 Trusted Process
| createObject(len,type) obj_id | Create data object (type=shared/pipe). |
@@ 5.3.2 Trustlet
Scheduling. The guest OS initiates the scheduling of trustlets via the invokeTrustlet() call. Hacher runtime enforces a run-to-completion execution model, i.e., Hacher executes a trustlet until completion for efficient execution [163], and returns the result to the guest OS. The scheduler also handles function chaining of the co-located trustlets (§ 5.4).
@@ 5.4 Confidential Network Architecture
Function chaining. Chaining multiple functions is a common pattern in serverless computing [35, 142, 157]. Hacher’s data object enables efficient function chaining through shared memory [128, 119]. Specifically, when chained functions are scheduled on the same host, the trusted monitor creates a special data object and grants write permission to the first function and read permission to the second function. If the functions can not be co-located, Hacher falls back to the normal network path: it returns the result of the first function to the serverless orchestrator, and then the orchestrator invokes the second function.
@@ 5.5 Differential Attestation
In confidential serverless computing, besides attesting the underlying platform, users need to verify the serverless runtime, the deployed function, and any inputs and outputs to validate the correct execution. This process can severely impact the function’s startup and execution latency. To alleviate this issue, Hacher introduces a differential attestation protocol to reduce the latency of verifying serverless function execution, which requires attesting the platform, runtime, function, and I/O. The key idea is to retain previously calculated measurements and compose a cumulative attestation report by measuring only the mutable parts at runtime for the function execution.
@@
Figure 7 illustrates the workflow of Hacher’s differential attestation protocol. It consists of two main phases: the initialization and the function invocation phase.
@@
The initialization phase begins with the function provider sharing the public function key with the user, which is used to encrypt the user requests. Then, during the Hacher initialization phase, the function provider establishes a secure connection with the Hacher instance to share the function’s private key and the function’s policy that specifies the measurements of zygotes and functions permitted for use. The establishment of the secure connection relies on the attestation of the trusted monitor, whose trust is rooted in the ASP [14].
@@
Precisely, the function provider first sends a nonce to the trusted monitor to prevent replay attacks. In turn, the attestation service generates a Diffie–Hellman (DH) key pair and retrieves the attestation report from the ASP, which includes the measurement of the trusted monitor, the provided nonce and the hash value of the DH public key as user data. The function provider receives and verifies the report and then sends the function private key and the policy to the Hacher instance. Hacher uses the function’s private key to decrypt user requests and the policy to attest the trusted processes.
@@
The function invocation phase includes two scenarios depending on the Hacher’s state, namely cold and warm start (Figure 4). In the cold start case, the attestation service verifies the integrity of the trusted process by comparing the measurements of its components against the values provided by the function provider’s policy during the initialization phase. Beyond this point, subsequent invocations of the same function become a warm start. Hacher’s attestation service only has to measure the function input and output and can generate a report based on prior measurements of the monitor, the zygote, and the invoked function. When several serverless functions are chained, the extended attestation report includes all their measurements. This approach reduces the attestation time while ensuring verifiability for the entire data chain.
@@ 6 Security Analysis
| Read the Hacher’s memory | VMPL protection |
@@
6.1 Attack vectors & Hacher’s mitigations
Table 4 summarizes potential attack vectors and Hacher’s mitigations.
@@ 6.1 Attack vectors & Hacher’s mitigations
Further, an attacker could load arbitrary zygotes and functions. However, Hacher’s attestation service blocks this process if there is a measurement discrepancy from the values in the function provider’s policy. Replay attacks by replacing user input are also detected as users can validate the integrity of the input based on the content of the attestation report. Lastly, if an attacker intercepts a file request from a trustlet to the nested filesystem (§ 5.2) and returns a compromised file, the LibOS verifies the file’s measurement against its expected value in Hacher’s manifest, rejecting any tampered files.
@@
Attempts to compromise Hacher by launching a tampered trusted monitor, bypassing it altogether, or directly booting Hacher in a standard VM are detected during Hacher’s remote attestation workflow. Attestation reports include invalid measurement values if altered software is used. Attempts to deploy Hacher as a normal VM can be identified, as, in this case, Hacher fails to produce a valid report from the ASP. Even crafted reports are detected, since only the CPU vendor’s key, securely managed by the ASP, can sign genuine attestation certificates. Lastly, establishing secure connections (e.g., TLS) protects against network communication manipulation by ensuring the confidentiality and integrity of network packets.
Co-located serverless functions. Hacher prevents attacks from co-located functions. An attacker might create a serverless function that exploits a runtime vulnerability aiming to access memory belonging to another trustlet. However, each trustlet runs in ring3 and has its own page tables managed by the trusted monitor, which prohibits access to other trustlet’s memory. Further, to prevent reuse-based attacks [162], Hacher’s process runtime recreates trustlets when receiving a request from a different user, resetting the potentially compromised residual state due to a previous invocation.
@@ 6.2 Formal verification of Protocols
We formally verify Hacher’s differential attestation protocol, as depicted in Figure 7, using the Tamarin Prover [103, 49] under the Dolev-Yao [50] attacker model. The verification considers two key properties for Hacher’s differential attestation protocol: (i) secrecy, secrets (e.g., encrypted data, private keys) remain undisclosed unless explicitly compromised; and (ii) authenticity, guaranteeing that attested results correspond to correct function execution with validated inputs. We model the steps specific to Hacher in detail.
@@
In this phase we also included an unrestricted machine_init to be able to set up an infinite number of machines and thus infinite potential Hacher instances. This, together with the way we modeled the function provider in the previous phase, allows us to model an unbounded number of Hacher instances, machines, functions, providers, and users all while ensuring they are unable to interfere to violate our desired properties.
@@
Lemmas. We use to denote that action fact a occurred at time . Leveraging these action facts, we verify the following lemmas for Hacher’s differential attestation protocol:
@@ 7 Implementation
Trusted monitor. We implement our prototype of Hacher for Linux environments. We base Hacher’s trusted monitor on COCONUT-SVSM [37]. COCONUT-SVSM is a service module that operates at VMPL-0 and aims to provide services (e.g., vTPM [111]) to a guest running at VMPL-2. We extend COCONUT-SVSM to support trusted processes in VMPL-1 and incorporate Hacher’s differential attestation service. We base our LibOS for trusted processes on Gramine [26, 38] and implement a backend to enable its execution at VMPL-1. Hacher’s current prototype uses a patched Linux version v6.8 (host OS) and v6.5 (guest OS), configured to support COCONUT-SVSM.
@@
LibOS. Our LibOS for trusted processes uses Gramine [26, 38] with a backend for VMPL-1 execution. Gramine defines a Platform Adaptation Layer (PAL) host ABI [39] to delegate specific operations to the host. We implement our PAL as a shim layer for our trusted monitor, which allows Hacher to run the Python runtime as a trusted process. Its core functionalities include memory allocation and external file access, and it leverages Hacher’s communication protocol to perform its PAL ABI operations requests.
@@
Attestation service. Hacher leverages COCONUT-SVSM’s get_regular_report() to retrieve a signed AMD SEV-SNP report and, then, stores it securely in the trusted monitor’s memory. To calculate the measurements of individual trusted processes components (e.g., zygote image, trustlet), Hacher uses the SHA-512 algorithm (same as the SEV-SNP’s measurement algorithm [14]). These measurements are cached in the zygote and trustlet contexts for faster retrieval. The monitor invokes the attestation functions internally when fetching a zygote or creating a trustlet, validating them against the function provider’s policy by comparing the calculated with the provided measurements. Finally, on a function invocation, the measurements of the used software components and the function input and output are appended to the trusted monitor’s attestation report, which is then signed using the private key provided by the function provider and forwarded to the user for verification of the entire execution.
Guest OS components. Hacher includes a kernel module that allows a guest application to make monitor calls. In addition, we implement a Python library that interacts with the trusted monitor through this kernel module. We also extend the SeBS benchmark suite [33] to support running serverless functions with Hacher using this Python library.

### v3

@@ 5.1 Trusted Monitor
Monitor system calls. A serverless framework, running within the guest, interacts with the trusted monitor through the defined monitor system calls (Table 2). These calls facilitate the execution of serverless functions, including the creation of zygotes/trustlets, function invocation, performing attestation, and loading the provider’s policy. Additionally, the trusted monitor defines an interface between trusted processes for the LibOS to request backend operations (§ 5.2). All interfaces are kept small and concise to eliminate the risk of possible attack vectors introduced by internal oversight.
@@
Attestation service. The trusted monitor includes an attestation service that generates an attestation report for remote users to verify the integrity of Wallet and its function execution. On function execution, the attestation service calculates the measurement of function components following Wallet’s differential attestation principles (§ 5.5) to optimize the measurement process by caching previously computed results. The root-of-trust of Wallet relies on CVM’s attestation mechanism [14, 74], whose attestation report includes the measurement of the initial guest state and memory. The attestation service integrates this report into its final result.
Performance consideration. The trusted monitor is a central part of Wallet’s operations. To ensure performant operations with multiple functions, the monitor allows each I/O operation to run simultaneously for multiple trustlets, without compromising its security properties. For memory accesses, the monitor is only called when (de)allocating memory. It manipulates the trustlets’ page table and sets the correct VMPL levels for the new pages, an operation which can also be parallelized.
@@ 5.2 LibOS Architecture
Wallet’s LibOS [23, 121, 31, 54, 91, 27] integrates a function runtime as well as a shim layer and filesystem with a nested namespace for efficient function execution.
@@
Filesystem. The LibOS provides an in-memory file system backed by a zygote, at the cost of a bloated image. Removing non-commonly-used files reduces image size but compromises serverless function flexibility and user experience. As a solution, Wallet employs a nested namespace [149] that integrates an embedded with an external filesystem. From the application’s perspective, all files appear at the same mount point (e.g., ‘/’). Internally, the LibOS first checks the embedded filesystem; if a file is missing, it retrieves it from the guest OS’s filesystem via the monitor. To ensure file integrity, the shim layer verifies files from the guest OS using provided measurements specified in a manifest [68, 153] embedded in the zygote image and aborts operations in case of a mismatch.
Data-centric I/O. For inter-function communication, the LibOS provides a data object abstraction to the trusted processes. This abstraction allows a function to create, get, and update data objects via the API shown in Table 3. Unlike conventional CVM networking that requires traversing entire I/O stacks with multiple context switches and encryption overhead [30, 92, 112, 62], Wallet employs a shared memory-based local object store [131, 161, 86, 122, 120, 29], enabling faster communication. This is especially effective for Wallet as it avoids the costly CVM network overhead (§ 3.3). When the serverless orchestrator fails to co-locate functions, the trusted monitor follows a hybrid I/O approach [101, 161] and resorts to the standard CVM I/O primitives via the guest OS, similar to the external file system access described above. Importantly, this mechanism ensures transparent communication between functions, regardless of whether they are co-located or distributed. § 5.4 details Wallet’s communication mechanism.
@@ 5.3 Trusted Process
| createObject(len,type) obj_id | Create data object. |
@@ 5.3.2 Trustlet
Scheduling. The guest OS initiates the scheduling of trustlets via the invokeTrustlet() call. Wallet runtime enforces a run-to-completion execution model, i.e., Wallet executes a trustlet until completion for efficient execution [168], and returns the result to the guest OS. This fits for serverless workloads, which are most of the time short-lived. In case a function performs I/O tasks, the trustlet returns control back to the monitor and a different trustlet can be scheduled, even before the original one has finished its execution. On top of that, the scheduler is also responsible for handling the function chaining of the co-located trustlets (§ 5.4).
@@ 5.4 Confidential Network Architecture
Function chaining. Chaining multiple functions is a common pattern in serverless computing [36, 147, 162]. Wallet’s data object enables efficient function chaining through shared memory [131, 122]. Specifically, when chained functions are scheduled on the same host, the trusted monitor creates a special data object and grants write permission to the first function and read permission to the second function. If the functions can not be co-located, Wallet falls back to the normal network path: it returns the result of the first function to the serverless orchestrator, and then the orchestrator invokes the second function.
Note that, these chains are not limited to two functions and could potentially result in a circular chain, leading to the functions not producing a result. The current Wallet prototype does not consider this case. However, circular chains can be detected and prevented by the monitor in a future version. or Wallet’s policies could be extended to explicitly disallow it.
External network I/O. To facilitate network communications in Wallet across the CVM boundaries, additional steps are required. The monitor delegates the external network I/O handling to the guest OS. In practice, the monitor should encrypt the network data and provide it to the serverless framework for forwarding. Then, the serverless framework can send this data to the desired destination and get a response.
@@ 5.5 Differential Attestation
In confidential serverless computing, besides attesting the underlying platform, users need to verify the serverless runtime, the deployed function, and any inputs and outputs to validate the correct execution. This process can severely impact the function’s startup and execution latency. To alleviate this issue, Wallet introduces a differential attestation protocol to reduce the latency of verifying serverless function execution, which requires attesting the platform, runtime, function, and I/O. The key idea is to retain previously calculated measurements and compose a cumulative attestation report by measuring only the mutable parts at runtime for the function execution.
Security guarantees. Wallet’s differential attestation ensures the integrity of the monitor, trustlet, zygote, function, its input, and output. It relies on the correctness of Wallet, in particular on Wallet’s monitor, which is the only permanently immutable component of the system. All other parts of the differential attestation are at one point mutable, since they are fetched and instantiated by the monitor. After their initialization, all parts that constitute a trustlet become immutable with data protected with CoW. This enables the monitor to cache the hashes of these components for future use.
@@
Attestation flow. Figure 7 illustrates the workflow of Wallet’s differential attestation protocol. It consists of two main phases: the initialization and the function invocation phase.
@@
The initialization phase begins with the function provider sharing the public function key with the user, which is used to encrypt the user requests. Then, during the Wallet initialization phase, the function provider establishes a secure connection with the Wallet instance to share the function’s private key and the function’s policy that specifies the measurements of zygotes and functions permitted for use. The establishment of the secure connection relies on the attestation of the trusted monitor, whose trust is rooted in the ASP [14].
@@
Precisely, the function provider first sends a nonce to the trusted monitor to prevent replay attacks. In turn, the attestation service generates a Diffie–Hellman (DH) key pair and retrieves the attestation report from the ASP, which includes the measurement of the trusted monitor, the provided nonce and the hash value of the DH public key as user data. The function provider receives and verifies the report and then sends the function private key and the policy to the Wallet instance. Wallet uses the function’s private key to decrypt user requests and the policy to attest the trusted processes.
Importantly, before providing private or security-sensitive data as a function input, a user must first perform a test function call, retrieve the attestation report, and verify that the Wallet instance is at the expected state and the correct toolchain is present. After this point, the user has gained trust in Wallet and can proceed with their function invocations.
@@
The function invocation phase includes two scenarios depending on the Wallet’s state, namely cold and warm start (Figure 4). In the cold start case, the attestation service verifies the integrity of the trusted process by comparing the measurements of its components against the values provided by the function provider’s policy during the initialization phase. Beyond this point, subsequent invocations of the same function become a warm start. Wallet’s attestation service only has to measure the function input and output and can generate a report based on prior measurements of the monitor, the zygote, and the invoked function. When several serverless functions are chained, the extended attestation report includes all their measurements. This approach reduces the attestation time while ensuring verifiability for the entire data chain.
@@ 6 Security Analysis
| Read the Wallet’s memory | VMPL protection |
@@
6.1 Attack vectors & Wallet’s mitigations
Table 4 summarizes potential attack vectors and Wallet’s mitigations.
@@ 6.1 Attack vectors & Wallet’s mitigations
Further, an attacker could load arbitrary zygotes and functions. However, Wallet’s attestation service blocks this process if there is a measurement discrepancy from the values in the function provider’s policy. Replay attacks by replacing user input are also detected as users can validate the integrity of the input based on the content of the attestation report. Lastly, if an attacker intercepts a file request from a trustlet to the nested filesystem (§ 5.2) and returns a compromised file, the LibOS verifies the file’s measurement against its expected value in Wallet’s manifest, rejecting any tampered files.
@@
Attempts to compromise Wallet by launching a tampered trusted monitor, bypassing it altogether, or directly booting Wallet in a standard VM are detected during Wallet’s remote attestation workflow. Attestation reports include invalid measurement values if altered software is used. Attempts to deploy Wallet as a normal VM can be identified, as, in this case, Wallet fails to produce a valid report from the ASP. Even crafted reports are detected, since only the CPU vendor’s key, securely managed by the ASP, can sign genuine attestation certificates. Lastly, establishing secure connections (e.g., TLS) protects against network communication manipulation by ensuring the confidentiality and integrity of network packets.
Co-located serverless functions. Wallet prevents attacks from co-located functions. An attacker might create a serverless function that exploits a runtime vulnerability aiming to access memory belonging to another trustlet. However, each trustlet runs in ring3 and has its own page tables managed by the trusted monitor, which prohibits access to other trustlet’s memory. Further, to prevent reuse-based attacks [167], Wallet’s process runtime recreates trustlets when receiving a request from a different user, resetting the potentially compromised residual state due to a previous invocation.
@@ 6.2 Formal verification of Protocols
We formally verify Wallet’s differential attestation protocol, as depicted in Figure 7, using the Tamarin Prover [105, 50] under the Dolev-Yao [51] attacker model. The verification considers two key properties for Wallet’s differential attestation protocol: (i) secrecy, secrets (e.g., encrypted data, private keys) remain undisclosed unless explicitly compromised; and (ii) authenticity, guaranteeing that attested results correspond to correct function execution with validated inputs. We model the steps specific to Wallet in detail.
@@
In this phase we also included an unrestricted machine_init to be able to set up an infinite number of machines and thus infinite potential Wallet instances. This, together with the way we modeled the function provider in the previous phase, allows us to model an unbounded number of Wallet instances, machines, functions, providers, and users all while ensuring they are unable to interfere to violate our desired properties.
@@
Lemmas. We use to denote that action fact a occurred at time . Leveraging these action facts, we verify the following lemmas for Wallet’s differential attestation protocol:
@@ 7 Implementation
Trusted monitor. We implement our prototype of Wallet for Linux environments. We base Wallet’s trusted monitor on COCONUT-SVSM [38]. COCONUT-SVSM is a service module that operates at VMPL-0 and aims to provide services (e.g., vTPM [114]) to a guest running at VMPL-2. We extend COCONUT-SVSM to support trusted processes in VMPL-1 and incorporate Wallet’s differential attestation service. We base our LibOS for trusted processes on Gramine [27, 39] and implement a backend to enable its execution at VMPL-1. Wallet’s current prototype uses a patched Linux version v6.8 (host OS) and v6.5 (guest OS), configured to support COCONUT-SVSM.
@@
LibOS. Our LibOS for trusted processes uses Gramine [27, 39] with a backend for VMPL-1 execution. Gramine defines a Platform Adaptation Layer (PAL) host ABI [40] to delegate specific operations to the host. We implement our PAL as a shim layer for our trusted monitor, which allows Wallet to run the Python runtime as a trusted process. Its core functionalities include memory allocation and external file access, and it leverages Wallet’s communication protocol to perform its PAL ABI operations requests.
@@
Attestation service. Wallet leverages COCONUT-SVSM’s get_regular_report() to retrieve a signed AMD SEV-SNP report and, then, stores it securely in the trusted monitor’s memory. To calculate the measurements of individual trusted processes components (e.g., zygote image, trustlet), Wallet uses the SHA-512 algorithm (same as the SEV-SNP’s measurement algorithm [14]). These measurements are cached in the zygote and trustlet contexts for faster retrieval. The monitor invokes the attestation functions internally when fetching a zygote or creating a trustlet, validating them against the function provider’s policy by comparing the calculated with the provided measurements. Finally, on a function invocation, the measurements of the used software components and the function input and output are appended to the trusted monitor’s attestation report, which is then signed using the private key provided by the function provider and forwarded to the user for verification of the entire execution.
Guest OS components. Wallet includes a kernel module that allows a guest application to make monitor calls. In addition, we implement a Python library that interacts with the trusted monitor through this kernel module. We also extend the SeBS benchmark suite [34] to support running serverless functions with Wallet using this Python library.